OpenStudy (anonymous):

convert the polar point (4 pi) into rectangular

OpenStudy (anonymous):

I got -4,0, please tell me if I am right

OpenStudy (turingtest):

if you mean the point \((4,\pi)\) then yep. you're right

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

yep confirmed

OpenStudy (anonymous):

@Posideon can u look at this? and @TuringTest : What type of discontinuity does the function y = (x^2 - 5x) / (x^2 + 5x) have at x = 0 ?

OpenStudy (anonymous):

take the lim as x goes to 0.

OpenStudy (anonymous):

I got Removable discontinuity at x = 0

OpenStudy (anonymous):

if you get a finite value then thats a removable discontinuity i think

OpenStudy (anonymous):

you should graph it to check

OpenStudy (anonymous):

can u graph? i dont have it

OpenStudy (anonymous):

but thats sounds good as the limit is -1. :D

OpenStudy (anonymous):

wolfram alpha search that on google